# DarkAI: Reconstructing the density, velocity and tidal field of dark matter from DESI-like bright galaxy sample

### v1

@@ Abstract
Reconstructing the dark matter density, velocity, and tidal (MTV) fields from galaxy surveys is essential for advancing our understanding of the large-scale structure of the Universe. In this work, we present a machine learning-based framework using a UNet convolutional neural network to reconstruct the MTV fields from mock samples of the DESI bright galaxy survey within the redshift range $0.40.1<z<$ 0.4. Our approach accounts for realistic observational effects, including geometric selection, flux-limited data, and redshift space distortion (RSD) effects, thereby improving the fidelity of the reconstructed fields. Testing on mock galaxy catalogs generated from the Jiutian N-body simulation, our method achieves significant accuracy level. The reconstructed density field exhibits strong consistency with the true field, effectively eliminating most RSD effects and achieving a cross-correlation power spectrum coefficient greater than 0.985 on scales with $k<0.1\,h\,\mathrm{Mpc}^{-1}$. The velocity field reconstruction accurately captures large-scale coherent flows and small-scale turbulent features, exhibiting slopes of the grid-to-grid relationship close to unity and a scatter below $\sim$100 $\mathrm{km}\,\mathrm{s}^{-1}$. Additionally, the tidal field is reconstructed without bias, successfully recovering the features of the large-scale cosmic web, including clusters, filaments, sheets, and voids. Our results confirm that the proposed framework effectively captures the large-scale distribution and dynamics of dark matter while addressing key systematic challenges. These advancements provide a reliable and robust tool for analyzing current and future galaxy surveys, paving the way for new insights into cosmic structure formation and evolution.
Large-scale structure of the universe(902) — Dark matter distribution(356) — Redshift surveys(1378)—Convolutional neural networks(1938)
@@ 1 Introduction
Understanding the distribution and dynamics of matter in the Universe provides critical insights into gravitational interactions and cosmic flows. This knowledge is essential for exploring the processes that drive the formation and evolution of cosmic structures and galaxies. To achieve this, galaxy redshift surveys, such as the 2-degree Field Galaxy Redshift Survey (Colless et al., 2001) and the Sloan Digital Sky Survey (SDSS) (York et al., 2000), offer an opportunity to probe the large-scale mass distribution of the Universe. In the coming decade, next-generation galaxy surveys, including the Dark Energy Spectroscopic Instrument (DESI) (DESI Collaboration et al., 2016a, b), the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019), EUCLID (Laureijs et al., 2011), the Nancy Grace Roman Space Telescope (Akeson et al., 2019), and the Chinese Space Station Telescope (CSST) (Zhan, 2011; Gong et al., 2019), are poised to map an unprecedented volume of the Universe with different redshift measurement accuracy. These surveys will enable transformative advances in our understanding of cosmic structure and evolution, emphasizing the importance of developing robust and efficient methods to reconstruct the three-dimensional dark-matter density, velocity, and tidal (MTV) field from the vast galaxy survey datasets.
@@
Reconstructing the MTV field from observational data is essential for addressing key scientific questions. Constrained simulations, such as those in the Local Universe (e.g., Wang et al., 2016; Tweed et al., 2017), require accurate initial conditions derived from the density field to reproduce observed structures. Local non-Gaussianity in the cosmic web can be quantified through the MTV field, offering insights into primordial fluctuations (Desjacques & Seljak, 2010; Smith et al., 2018; Münchmeyer et al., 2019). Filament spin studies, such as alignment of the galaxy angular momentum with filaments (e.g., Tempel & Libeskind, 2013; Wang et al., 2021), rely on the reconstructed tidal field for precise measurements. Filament lensing studies, mapping dark matter in filaments (e.g., HyeongHan et al., 2024), are similarly dependent on accurate reconstructions. Moreover, correcting supernova distance measurements for mass density and peculiar velocities (Cooray & Caldwell, 2006; Hui & Greene, 2006; Davis et al., 2011; Huang et al., 2024) and mapping baryon distributions using the velocity field (e.g., Dolag & Sunyaev, 2013; Yu & Zhu, 2019) underscore the broad utility of MTV reconstructions.
@@
Reconstructing the cosmic density and velocity fields has been conducted earlier from several redshift surveys (Fisher et al., 1995; Zaroubi et al., 1995; Schmoldt et al., 1999; Mathis et al., 2002; Erdoǧdu et al., 2004; Wang et al., 2009, 2013). Early efforts (Fisher et al., 1995; Zaroubi et al., 1995) relied on smoothing and normalizing galaxy data to approximate large-scale density fields, with methods such as Wiener reconstruction minimizing errors through linear combinations of galaxy density values. Other approaches (Wang et al., 2009, 2012, 2013) leveraged dark-matter halos or galaxy groups, assuming a linear bias between halo and mass density fields. However, these methods are quite complicated. The assumption of a scale-independent bias parameter lacks robust physical justification and is often motivated by simplicity, reducing its reliability in capturing complex relationships (e.g., Coles & Erdogdu, 2007; Simon & Hilbert, 2018). Moreover, linear theory assumptions, which posit a simple proportionality between gravity and velocity fields, fail in the nonlinear clustering regime due to phenomena like shell crossing, where the one-to-one correspondence between density and velocity fields breaks down. Additionally, redshift space distortions (RSDs) (Sargent & Turner, 1977; Davis & Peebles, 1983; Kaiser, 1987; Hamilton, 1992), influenced by the Kaiser effect on large scales (Kaiser, 1987) and the Finger-of-God effect on small scales (Jackson, 1972; Tully & Fisher, 1978), intertwine complex velocity and density field distortions, further challenging accurate reconstructions. Methods based on linear theory or the Zeldovich approximation are particularly limited in high-density regions, emphasizing the need for innovative strategies to fully leverage redshift survey data in mapping the Universe’s large-scale structure.
@@
Several studies have employed machine learning (ML) techniques to reconstruct dark matter density and velocity fields, demonstrating significant potential for advancing the field. A UNet-based neural network was used to reconstruct peculiar velocity fields from redshift-space distributions of dark matter halos, achieving high accuracy at multiple scales(Wu et al., 2021, 2023; Qin et al., 2023; Wang et al., 2024; Xiao et al., 2024). However, these approaches face notable limitations. First, these approaches often operate within idealized simulation environments, with limitations such as fixed box volumes, periodic boundary conditions, and the exclusion of observational systematics like geometric selection and flux-limited effects. Second, reconstruction based on the halo distributions compared to the galaxies requires additional bias modeling to bridge the gap between galaxy and dark matter. Addressing these limitations is essential for extending the ML-based reconstruction to real-world cosmological observations.
@@
In this paper, we aim to develop a robust ML-based framework to reconstruct the MTV fields from DESI-like bright galaxy samples. In Wang et al. (2024, hereafter Wang24), the first paper in this series, we present a deep-learning technique for reconstructing the dark-matter density field from the redshift-space distribution of dark-matter halos. The reconstruction was clearly proved to be able to correct for redshift-space distortions and is unaffected by the different cosmologies between the training and test samples. It is also tested that the application of the UNet-reconstructed density field can accurately infer the velocity & tidal field accurately, providing better results compared to the traditional approach based on the linear bias model. Here we improve upon the reconstruction method of Wang24 by addressing realistic observation challenges, including accounting for geometric selection, varying redshift distributions, and flux-limited observational biases. By incorporating systematic effects into the ML model, we aim to enhance the fidelity of the reconstructed fields and expand their applicability to real survey data. This study concentrates solely on reconstruction using the distribution of galaxies. We will address reconstruction involving groups or halos derived from observational data in a future probe.
@@
To support the DESI galaxy survey, we generate mock galaxy samples along with their corresponding dark matter density and velocity fields within a DESI-like light cone. These mock data sets are essential for studying the mapping between galaxies and the underlying large-scale structure of the universe, as well as for testing and validating cosmological analysis pipelines. The mock samples are constructed through a multi-step process that includes N-body simulations, halo and subhalo identification, galaxy population modeling, and the incorporation of survey selection effects. This section outlines the methodology utilized to generate these mock samples, based on the procedures described in Gu et al. (2024).
@@
The foundation of our mock galaxy samples is the Jiutian N-body simulation, which was run at the High-Performance Computing Center at Kunshan using L-GADGET, a memory-optimized version of the GADGET2 code (Springel, 2005). The Jiutian simulation describes the distribution of $6144^{3}$ dark-matter particles in a periodic box of $1000\>h^{-1}{\rm{Mpc}}$, with a particle mass of $3.723\times 10^{8}h^{-1}M_{\odot}$. The cosmological parameters were set to be consistent with the Planck2018 results (Planck Collaboration et al., 2020): $\Omega_{m}=0.3111$, $\Omega_{\Lambda}=0.6889$, $\Omega_{b}=0.049$, $\sigma_{8}=0.8102$, and $n_{s}=0.9665$.
@@
Dark matter haloes are identified using the friends-of-friends (FOF) algorithm (Davis et al., 1985) with a linking length of 0.2 times the mean interparticle separation. Subhaloes and their evolutionary histories are further processed using the Hierarchical Bound-Tracing (HBT+) algorithm (Han et al., 2012, 2018), which operates in the time domain to track the evolution of each halo throughout the simulation. The minimum number of particles in a subhalo is set to 20, corresponding to a minimum halo mass of approximately $7.5\times 10^{9}h^{-1}M_{\odot}$. The HBT+ code provides high-quality subhalo catalogues and a robust merger tree by following the most-bound particle even if a subhalo is no longer resolved.
@@
We populate the dark matter halos in our simulation with mock galaxies using the Subhalo Abundance Matching (SHAM) method. This technique establishes a connection between the properties of galaxies and their corresponding subhaloes by assuming a monotonic relationship between these properties (Kravtsov et al., 2004; Vale & Ostriker, 2004; Conroy et al., 2006; Yang et al., 2012; Wechsler & Tinker, 2018). The positions and velocities of galaxies are assigned by directly inheriting the positions and velocities of their host subhaloes from the N-body simulations. The luminosities are simulated by matching the $z$-band cumulative galaxy luminosity functions measured from the DESI One-percent survey (DESI Collaboration et al., 2024a) with the cumulative subhalo mass functions derived from the peak mass of all the subhaloes. To reflect the luminosity-subhalo mass relation more realistically, a scatter in the $z$-band luminosity, $\sigma_{\mathrm{log}}(L_{z})=0.15$ dex (e.g. Yang et al., 2008), is added to each galaxy.
@@
Figure 2: Normalized redshift distribution. Blue solid and red dashed lines correspond to the DESI Y1 catalog and our mock DESI sample, respectively.
We construct mock DESI survey and its corresponding dark matter particle distribution using the following steps:
@@
    For galaxies, we assign each a redshift, $z_{\rm obs}$, including the effects of redshift-space distortion, calculated as:
@@
    where $z_{\rm cos}$ is the cosmological redshift from Hubble expansion; $z_{\rm pec}$ is the Doppler contribution due to the galaxy’s line-of-sight peculiar velocity; $v_{\rm pec}$, $c$ is the speed of light. For dark matter particles, we retain them in real space.
@@
    For galaxies, we assign a $z$-band apparent magnitude, $m_{z}$, based on their distance and luminosity, incorporating a $K$-correction described as:
@@
In this study, we focus on the BGS sample with a $z$-band magnitude limit cut $m_{z}<19.0$ and within the redshift range of $0.40.1<z<$ 0.4, which is consistent with the DESI DR1 BGS (DESI Collaboration et al., 2024b). Figure 2 shows the normalized redshift distributions for our mock DESI sample (blue solid curve). For comparison, we also include the redshift distribution derived from the DESI Y1 catalogue (black dashed line). The redshift distribution of our mock sample generally follows the observed trend. Since the DESI Y1 catalogue covers a relatively small fraction of the sky, a slight difference between the two distributions is expected due to cosmic variance. In addition, DESI will cover two sky regions: a larger region in the Northern Galactic Cap (NGC) and a smaller region in the Southern Galactic Cap (SGC). To use a more contiguous region, we apply our reconstruction on the mainland of NGC, as shown in the red contour of Figure 1. The final sky coverage used in our reconstruction is about 9625.23 deg2. In this study, we construct a set of 12 mock samples by randomly rotating and shifting the stacked simulation boxes, which are used for training and testing our network.
@@
To construct the grid-based fields for dark matter and galaxies within the DESI-like survey volume, we proceed as follows. First, we embed the survey volume within a periodic cubic box with a side length of $1564\,h^{-1}\,\mathrm{Mpc}$. To maximize the utilization of the box volume, we apply a constant angular rotation in the right ascension (R.A.) direction. The 3D positions $(x,y,z)$ of the objects are then calculated by
@@
where $r$ is the line-of-sight comoving distance, $\beta$ is the declination (Dec.), $\alpha$ is R.A., and $\alpha_{\mathrm{rot}}=58^{\circ}$ is the applied rotation angle in this study.
We use the cloud-in-cell (CIC) scheme (Hockney & Eastwood, 1981) to compute the overdensity field, $\delta$, on a grid of $512^{3}$ grids for both dark matter and galaxy distributions. The velocity field, $\boldsymbol{v}$, is also estimated using the CIC method and normalized by the number of dark matter particles in each cell of the grid. To suppress shot noise, we smooth both $\delta$ and $\boldsymbol{v}$ fields using a TensorFlow convolutional kernel with a kernel size of 3. Figure 3 shows the projected field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. From left to right, the panels correspond to the galaxy overdensity, dark matter density, and velocity magnitude $|\boldsymbol{v}|$, respectively. The correlated structures among the three fields are clearly visible.

### v2

# DarkAI: Reconstructing the Density, Velocity and Tidal Fields of Dark Matter from a DESI-like Bright Galaxy Sample

@@ Abstract
Reconstructing the mass density, velocity, and tidal (MTV) fields of dark matter from galaxy surveys is essential for advancing our understanding of the LSS of the Universe. In this work, we present a machine learning-based framework using a UNet convolutional neural network to reconstruct the MTV fields from mock samples of the DESI bright galaxy survey within the redshift range $0.1<z<0.4$. Our approach accounts for realistic observational effects, including geometric selection, flux-limited data, and redshift space distortion (RSD) effects, thereby improving the fidelity of the reconstructed fields. Testing on mock galaxy catalogs generated from the Jiutian N-body simulation, our method achieves significant accuracy level. The reconstructed density field exhibits strong consistency with the true field, effectively eliminating most RSD effects and achieving a cross-correlation power spectrum coefficient greater than 0.985 on scales with $k<0.1\,h\,\mathrm{Mpc}^{-1}$. The velocity field reconstruction accurately captures large-scale coherent flows and small-scale turbulent features, exhibiting slopes of grid-to-grid relationships close to unity and scatter below $\sim$100 $\mathrm{km}\,\mathrm{s}^{-1}$. Additionally, the tidal field is reconstructed without bias, successfully recovering the features of the large-scale cosmic web, including clusters, filaments, sheets, and voids. Our results confirm that the proposed framework effectively captures the large-scale distribution and dynamics of dark matter while addressing key systematic challenges. These advancements provide a reliable and robust tool for analyzing current and future galaxy surveys, paving the way for new insights into cosmic structure formation and evolution.
LSS of the universe(902) — Dark matter distribution(356) — Redshift surveys(1378)—Convolutional neural networks(1938)
@@ 1 Introduction
Understanding the distribution and dynamics of matter in the Universe provides critical insights into gravitational interactions and cosmic flows. This knowledge is essential for exploring the processes that drive the formation and evolution of cosmic structures and galaxies. To achieve this, galaxy redshift surveys, such as the 2-degree-field Galaxy Redshift Survey (Colless et al., 2001) and the Sloan Digital Sky Survey (York et al., 2000), offer an opportunity to probe the large-scale mass distribution of the Universe. In the coming decade, next-generation galaxy surveys, including the Dark Energy Spectroscopic Instrument (DESI) (DESI Collaboration et al., 2016a, b), the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (Ivezić et al., 2019), EUCLID (Laureijs et al., 2011), the Nancy Grace Roman Space Telescope (Akeson et al., 2019), and the Chinese Space Station Telescope (e.g., Zhan, 2011; Gong et al., 2019; Shi et al., 2025), are poised to map an unprecedented volume of the Universe with different redshift measurement accuracies. These surveys will enable transformative advances in our understanding of cosmic structure and evolution, emphasizing the importance of developing robust and efficient methods for reconstructing the three-dimensional mass density, velocity, and tidal (MTV) fields of dark matter from the vast galaxy survey datasets.
@@
Reconstructing the MTV field from observational data is essential for addressing key scientific questions. Constrained simulations, such as those in the local Universe (e.g., Wang et al., 2016; Tweed et al., 2017), require accurate initial conditions derived from the density field to reproduce observed structures. Local non-Gaussianity in the cosmic web can be quantified through the MTV field, offering insights into primordial fluctuations (Desjacques & Seljak, 2010; Smith et al., 2018; Münchmeyer et al., 2019). Filament spin studies, such as the alignment of the galaxy angular momentum with filaments (e.g., Tempel & Libeskind, 2013; Wang et al., 2021), rely on the reconstructed tidal field for precise measurements. Filament lensing studies, mapping dark matter in filaments (e.g., HyeongHan et al., 2024), are similarly dependent on accurate reconstructions. Moreover, correcting supernova distance measurements for mass density and peculiar velocities (Cooray & Caldwell, 2006; Hui & Greene, 2006; Davis et al., 2011; Huang et al., 2024) and mapping baryon distributions using the velocity field (e.g., Dolag & Sunyaev, 2013; Yu & Zhu, 2019) underscore the broad utility of MTV reconstructions.
@@
Reconstructing the cosmic density and velocity fields has previously been conducted from several redshift surveys (Fisher et al., 1995; Zaroubi et al., 1995; Schmoldt et al., 1999; Mathis et al., 2002; Erdoǧdu et al., 2004; Wang et al., 2009, 2013). Early efforts (Fisher et al., 1995; Zaroubi et al., 1995) relied on smoothing and normalizing galaxy data to approximate large-scale density fields, with methods such as Wiener reconstruction minimizing errors through linear combinations of galaxy density values. Other approaches (Wang et al., 2009, 2012, 2013) leveraged dark matter halos or galaxy groups, assuming a linear bias between halo and mass density fields. However, these methods are quite complicated. The assumption of a scale-independent bias parameter lacks robust physical justification and is often motivated by simplicity, reducing its reliability in capturing complex relationships (e.g., Coles & Erdogdu, 2007; Simon & Hilbert, 2018). Moreover, linear theory assumptions, which posit a simple proportionality between gravity and velocity fields, fail in the nonlinear clustering regime, due to phenomena like shell crossing, where the one-to-one correspondence between density and velocity fields breaks down. Additionally, redshift space distortions (RSDs Sargent & Turner, 1977; Davis & Peebles, 1983; Kaiser, 1987; Hamilton, 1992), influenced by the Kaiser effect on large scales (Kaiser, 1987) and the Finger-of-God effect on small scales (Jackson, 1972; Tully & Fisher, 1978), intertwine complex velocity and density field distortions, further challenging accurate reconstructions. Methods based on linear theory or the Zel’dovich approximation are particularly limited in high-density regions (e.g., Shi et al., 2016, 2018), emphasizing the need for innovative strategies to fully leverage redshift survey data in mapping the Universe’s LSS (LSS).
@@
Several studies have employed machine learning (ML) techniques to reconstruct dark matter density or velocity fields, demonstrating significant potential for advancing the field (Wu et al., 2021, 2023; Chen et al., 2023; Qin et al., 2023; Wang et al., 2024b; Xiao et al., 2024). However, these approaches face notable limitations. First, these approaches often operate within idealized simulation environments, with limitations such as fixed box volumes, periodic boundary conditions, and the exclusion of observational systematics like geometric selection and flux-limited effects. Second, reconstruction based on the halo distributions compared to the galaxies requires additional bias modeling to bridge the gap between the galaxies and dark matter. Addressing these limitations is essential for extending the ML-based reconstruction to real-world cosmological observations.
@@
In this paper, we aim to develop a robust ML-based framework to reconstruct the MTV fields from a DESI-like bright galaxy survey (BGS). In Wang et al. (2024b, hereafter Wang24), the first paper in this series, we presented a deep learning technique for reconstructing the mass density field from the redshift-space distribution of dark matter halos. The reconstruction was clearly proved to be able to correct for RSDs and was unaffected by the different cosmologies between the training and test samples. It was also tested that the application of a UNet-reconstructed density field can accurately infer the velocity and tidal field, providing better results compared to a traditional approach based on a linear bias model. Here, we improve upon the reconstruction method of Wang24 by addressing realistic observation challenges, including accounting for geometric selection, varying redshift distributions, and flux-limited observational biases. By incorporating systematic effects into the ML model, we aim to enhance the fidelity of the reconstructed fields and expand their applicability to real survey data. This study concentrates solely on reconstruction using the distribution of galaxies. We will address reconstruction involving groups or halos derived from observational data in a future probe.
@@
To support the DESI galaxy survey, we generate mock galaxy samples along with their corresponding mass density and velocity fields of dark matter within a DESI-like light cone. These mock data sets are essential for studying the mapping between galaxies and the underlying LSS of the Universe, as well as for testing and validating cosmological analysis pipelines. The mock samples are constructed through a multistep process that includes N-body simulations, halo and subhalo identifications, galaxy population modeling, and the incorporation of survey selection effects. This section outlines the methodology utilized to generate these mock samples, based on the procedures described in Gu et al. (2024).
@@
The foundation of our mock galaxy samples is the Jiutian N-body simulation (Han et al., 2025), which was run at the High-Performance Computing Center at Kunshan using L-GADGET, a memory-optimized version of the GADGET2 code (Springel, 2005). The Jiutian simulation describes the distribution of $6144^{3}$ dark matter particles in a periodic box of $1000\>h^{-1}{\rm{Mpc}}$, with a particle mass of $3.723\times 10^{8}h^{-1}M_{\odot}$. The cosmological parameters were set to be consistent with the Planck2018 results (Planck Collaboration et al., 2020): $\Omega_{m}=0.3111$, $\Omega_{\Lambda}=0.6889$, $\Omega_{b}=0.049$, $\sigma_{8}=0.8102$, and $n_{s}=0.9665$.
@@
Dark matter halos are identified using the friends-of-friends algorithm (Davis et al., 1985) with a linking length of 0.2 times the mean interparticle separation. Subhalos and their evolutionary histories are further processed using the Hierarchical Bound-Tracing (HBT+) algorithm (Han et al., 2012, 2018), which operates in the time domain, to track the evolution of each halo throughout the simulation. The minimum number of particles in a subhalo is set to 20, corresponding to a minimum halo mass of approximately $7.5\times 10^{9}h^{-1}M_{\odot}$. The HBT+ code provides high-quality subhalo catalogs and a robust merger tree by following the most bound particle, even if a subhalo is no longer resolved.
@@
We populate the dark matter halos in our simulation with mock galaxies using the subhalo abundance matching method. This technique establishes a connection between the properties of galaxies and their corresponding subhalos, by assuming a monotonic relationship between these properties (Kravtsov et al., 2004; Vale & Ostriker, 2004; Conroy et al., 2006; Yang et al., 2012; Wechsler & Tinker, 2018). The positions and velocities of the galaxies are assigned by directly inheriting the positions and velocities of their host subhalos from the N-body simulations. The luminosities are simulated by matching the $z$-band cumulative galaxy luminosity functions measured from the DESI One-percent survey (DESI Collaboration et al., 2024a) with the cumulative subhalo mass functions derived from the peak mass of all the subhalos. To reflect the luminosity-subhalo mass relation more realistically, a scatter in the $z$-band luminosity, $\sigma_{\mathrm{log}}(L_{z})=0.15$ dex (e.g. Yang et al., 2008), is added to each galaxy.
@@
Figure 2: Normalized redshift distribution. The red dashed line shows the distribution for our mock DESI sample. The blue solid line shows the DESI galaxy sample, constructed by crossmatching photometric data from DESI LS DR9 with Year 1 spectroscopic observations.
Figure 3: Comparison of galaxy clustering between mock samples and DESI BGS measurements. The left panel displays the multipole moments of the 2PCF, while the right panel shows the corresponding power spectra. The lines represent results from the mock samples, while the symbols denote the DESI BGS measurements.
We construct a mock DESI survey and its corresponding dark matter particle distribution using the following steps:
@@
    For galaxies, we assign each a redshift, $z_{\rm obs}$, including the effects of RSD, calculated as:
@@
    where $z_{\rm cos}$ is the cosmological redshift from Hubble expansion, $z_{\rm pec}$ is the Doppler contribution due to the galaxy’s line-of-sight peculiar velocity, $v_{\rm pec}$, and $c$ is the speed of light. For dark matter particles, we retain them in real space.
@@
    For galaxies, we assign a $z$-band apparent magnitude, $m_{z}$, based on their distance and luminosity, incorporating a $K$-correction, described as:
@@
In this study, we focus on the BGS sample, with a $z$-band magnitude limit cut $m_{z}<19.0$ and within the redshift range of $0.1<z<0.4$, which is consistent with DESI DR1 BGS (DESI Collaboration et al., 2024b). Figure 2 shows the normalized redshift distributions for our mock DESI sample (red dashed line). For comparison, we also include the redshift distribution derived from the DESI galaxy sample (blue solid line), which is generated by Wang et al. (2024a). This sample is based on photometric imaging from the DESI Legacy Imaging Surveys (LS) DR9, with initial redshift estimates provided by the photometric redshifts for the LS catalog (Zhou et al., 2021). For galaxies with available spectra from DESI Year 1, redshifts are updated using fastspecfit measurements, adopting the recommended “best” redshift in cases of multiple observations. To match the selection criteria of BGS, we further restrict the sample to galaxies with $m_{z}<19.0$ and $0.1<z<0.4$. The redshift distribution of our mock sample generally follows the observed trend. In addition, DESI will cover two sky regions: a larger region in the Northern Galactic Cap (NGC) and a smaller region in the Southern Galactic Cap. To use a more contiguous region, we apply our reconstruction to the main region of the NGC, as shown in the red contour of Figure 1. The final sky coverage used in our reconstruction is about 9625.23 deg2. In this study, we construct a set of 12 mock samples by randomly rotating and shifting the stacked simulation boxes, which are used for training and testing our network.
To evaluate the consistency between our mock galaxy samples and the DESI observations, Figure 3 shows the comparison of the galaxy clustering. The left panels display the multipole moments of the two-point correlation function (2PCF), while the right panels show the corresponding power spectra. The lines represent the mock results, and the symbols denote the DESI results (DESI Collaboration et al., 2024c), measured from the BGS with $0.1<z<0.4$ and $M_{r}<-21.5$. The same selection criteria are applied to the mock sample to ensure consistency. Overall, both the 2PCF and power spectra exhibit reasonable agreement between the mock and DESI samples, indicating that the mock sample broadly reproduces the observed large-scale distributions. However, noticeable discrepancies remain, particularly in the monopole, primarily because we have not applied any clustering fitting in the mock sample relative to the observational data. We plan to incorporate such clustering fitting in future analyses, using the more robust DESI DR2 or DR3 data releases. Nevertheless, the current comparison demonstrates that our mock sample follows the DESI clustering distribution well and provides a reliable basis for testing the reconstruction framework and evaluating systematic effects.
@@
To construct the grid-based fields for dark matter and galaxies within the DESI-like survey volume, we proceed as follows. First, we embed the survey volume within a periodic cubic box with a side length of $1564\,h^{-1}\,\mathrm{Mpc}$. To maximize the utilization of the box volume, we apply a constant angular rotation in the R.A. direction. The 3D positions $(x,y,z)$ of the objects are then calculated by
@@
where $r$ is the line-of-sight comoving distance, $\beta$ is the declination (Dec.), $\alpha$ is the R.A., and $\alpha_{\mathrm{rot}}=58^{\circ}$ is the applied rotation angle in this study.
We use the cloud-in-cell (CIC) scheme (Hockney & Eastwood, 1981) to compute the overdensity field, $\delta$, on a grid of $512^{3}$ grids for both dark matter and galaxy distributions. The velocity field, $\boldsymbol{v}$, is also estimated using the CIC method and normalized by the number of dark matter particles in each cell of the grid. To suppress the shot noise, we smooth both the $\delta$ and $\boldsymbol{v}$ fields using a TensorFlow convolutional kernel, with a kernel size of 3. Figure 4 shows the projected field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. From left to right, the panels correspond to the galaxy overdensity, dark matter density, and velocity magnitude $|\boldsymbol{v}|$, respectively. The correlated structures among the three fields are clearly visible.
In this study, we focus on evaluating the reconstruction performance using mock DESI samples, accounting for geometric selection, flux limits, and RSDs. However, another systematic effect, namely the fiber assignment (FA), can also introduce biases into the reconstruction. We include a brief discussion of this effect in Appendix A. However, a comprehensive analysis of the FA effect and its correction is beyond the scope of this paper and will be addressed in future work.